3x - 2t = 4
2x -5 = -t

Found 2 solutions by addingup, stanbon:
Answer by addingup(3677)

3x - 2t = 4 (1)
--------------
2x - 5 = -t on this equation add t and 5 on both sides and you get:
2x + t = 5 now multiply all times 2:
4x + 2t = 10 add this equation to (1):
-------------------------------------------------
3x - 2t = 4
+
4x + 2t = 10
--------------
7x 0t = 14
7x = 14
x = 14/7 = 2
now that you know x = 2, solve for t:
3x - 2t = 4
3(2) - 2t = 4
6 - 2t = 4
-2t = -2
t = -2/-2 = 1
-------------------
Check, to check use the other equation:
2x - 5 = -t
2(2) - 5 = -1
4 - 5 = -1 Correct

Answer by stanbon(75887)

3x - 2t = 4
2x -5 = -t
-------
Rearrange::
3x-2t = 4
2x+t = 5
-------
Modify for elimination:
3x-2t = 4
4x+2t = 10
------
Add and solve for "x":
7x = 14
x = 2
Solve for "t"
t = 5-2x
t = 5-2*2
t = 1
-------------
Cheers,
Stan H.
-----------